OpenStudy (mendicant_bias):

I want to make sure that I'm correctly creating a linear combination of Heaviside functions from a given piecewise-defined function. Posted below in a moment.

OpenStudy (mendicant_bias):

My piecewise-defined function is: \[f(t) = \begin{cases}t, \ \ \ 0 \leq t<1\\0, \ \ \ t \ge1 \end{cases} \]

OpenStudy (mendicant_bias):

Using the general formula for a piecewise-defined function, I'm getting\[f(t)=t-t \ \mathcal{U}(t-1)\]and that's it.

OpenStudy (sidsiddhartha):

 graph should be this